OpenStudy (anonymous):

(5x-8)(4x+1) expand and simplify

OpenStudy (anonymous):

5x*4x=20x^2 -8*4x=32x 5x*1=5x -8*1=-8 20x^2+32x+5x-8 20x^2+37x-8

OpenStudy (anonymous):

wouldn't (-8*4x)= -32x

OpenStudy (anonymous):

Yes it would, my bad

OpenStudy (anonymous):

So then the -32x+5x would be -27x

OpenStudy (anonymous):

so 20x^2-27x+8, is your final value, just walk back through the steps to identify where easy came from.

OpenStudy (anonymous):

ohhh thank you

OpenStudy (anonymous):

each*

OpenStudy (anonymous):

is x^2 mean squared

OpenStudy (anonymous):

yes, it means x to the power of 2 so x^2 is squared x^3 is cubed and so on

OpenStudy (anonymous):

oh alright thank you :)

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

props also to paandaabear

OpenStudy (anonymous):

why is it 20^2 and not just 20 ?

OpenStudy (anonymous):

because quadratics are in ax^2+bx+c form x*x=x^2

OpenStudy (anonymous):

you can multiply out 5*4 to get 20 but there is an x term, which cant multiply but we use the notation as though it could, and since a number multiplied by itself is the same as squaring it, the same applies to x if we were to replace x with a value say x=3, then for x^2 we get 3^2 = 9

OpenStudy (anonymous):

ohh yeah okay thanks

OpenStudy (anonymous):

would (x-2)^2 be x^2-4x-4

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

okay and would (5x+4)(5x-4) be 10x-40x-16 ??

OpenStudy (anonymous):

no, its not 2*5x its 5x*5x , so = (5x*5x)-(5x*4)+(4*5x)-(4*4) = 25x^2-20x+20x-16 = 25x^2-16, since -20x+20x = 0

OpenStudy (anonymous):

5x*5x = 5^2 and x^2 = 25x^2, which is really 25 (x^2) 5*5x = 5^2 and x^1 = 25x

OpenStudy (anonymous):

ohh so whats the actual answer?

OpenStudy (anonymous):

for which? (5x+4)(5x-4)?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

= (5x*5x)-(5x*4)+(4*5x)-(4*4) = 25x^2-20x+20x-16 = 25x^2-16

OpenStudy (anonymous):

oh thank you soo much

OpenStudy (anonymous):

glad to help